# Engineering interaction potentials for stabilizing quantum quasicrystal phases

### v1

@@ VI Final Remarks
In this work, we systematically studied the necessary conditions for the stabilization of eight-, ten- and twelve-fold self-organized quantum quasicrystals. Unlike the classical scenario, we found that quantum quasicrystals with distinct rotational symmetries need pair interaction potentials with a different number of properly selected unstable wave vectors. In the case of the dodecagonal quasicrystal, a two negative minima potential is enough for the stabilization, as previously pointed out in the literature Heinonen et al. (2019). Meanwhile, the decagonal and octagonal quasicrystal structures need pair interaction potentials with at least three and four minima negative minima properly engineered to be stabilized, respectively. The greater stability of the dodecagonal quasicrystal with respect to other quasicrystalline structures considered resides in the geometric property of having six-fold rotational symmetry. Ultimately, this property is responsible for the existence of Fourier-mode triplets with wave vectors of equal modulus that adds up to zero. As occurs in hexagonal crystals, this feature is responsible for cubic potential energy contributions with a significant impact on decreasing the total energy cost of phases that share this symmetry, see Section III. Moreover, as can be observed in Fig. 1(f), Fig. 3(f) and Fig. 5(f), all quasicrystalline phases studied have a sector in which the superfluid fraction is significant, allowing us to conclude that these are actually super quasicrystal phases. These results were confirmed by extensive Gross-Pitaevskii computations, which produce equivalent configuration and superfluid fraction to those obtained with our variational spectral approach.
@@
We believe that the present results are relevant for the investigation of quasicrystal structures produced in cavity mediated interacting Bose-Einstein condensates Masalaeva et al. (2023); Mivehvar et al. (2021); Orsi et al. (2024); Mivehvar et al. (2019a). The design on demand of cavity mediated interactions have paved a way to investigate a variety of cutting-edge many-body phases. In this regard, such systems show the interesting opportunity to operate with lattice phononic modes, a flexible manner to test quasicrystals in a quantum framework. At the same time, the interplay between long-range magnetic and light-induced interactions in a Bose-Einstein condensateMishra et al. (2023) is another promising experimental setup that, in principle, should observe 8-QC, 10-QC and 12-QC patterns displaying a finite superfluid signal. Especially relevant is also the usage of dipolar atoms in cavitiesBonifacio et al. (2024b), allowing the implementation of many-body systems with highly tunable interactions capable of producing self-assembled quasicrystal clusters. Here it will be of key importance to inspect the stability of octagonal as well as decagonal phases.
@@ VII Acknowledgments
A.M.C. acknowledge UNIFI for financial support and hospitality. This work was supported by the European Union through the Next Generation EU funds through the Italian MUR National Recovery and Resilience Plan, Mission 4 Component 2–Investment 1.4–National Center for HPC, Big Data and Quantum Computing (CUP Grant No. B83C22002830001). F. C. and V. Z. acknowledge financial support from PNRR Ministero Università e Ricerca Project No. PE0000023-NQSTI funded by European Union-Next-Generation EU. The numerical solution of the GPE was performed using the XMDS2 software Dennis et al. (2013).

### v2

@@ VI Final Remarks
In this work, we systematically studied the necessary conditions for the stabilization of eight-, ten-, and twelve-fold self-organized quantum quasicrystals. Unlike the classical scenario, we found that quantum quasicrystals with distinct rotational symmetries require pair interaction potentials featuring different numbers of properly selected unstable wave vectors. In the case of the dodecagonal quasicrystal, a two-minima potential with negative values is sufficient for stabilization, as previously noted in the literature [heinonen2019quantum]. Meanwhile, the decagonal and octagonal quasicrystalline structures require pair interaction potentials with at least three and four properly engineered negative minima to be stabilized, respectively.
The smaller number of unstable wave vectors required to stabilize the dodecagonal quasicrystal ground-state indicates that this structure is more robust compared to the other cases considered. As discussed in Sec. III, this enhanced stability ultimately stems from the sixfold rotational symmetry of the pattern. This symmetry allows for the presence of Fourier-mode triplets with wave vectors of equal modulus forming an equilateral triangle whose vector sum is zero. As in hexagonal crystals, this property enables cubic Fourier amplitude contributions to the potential energy, which play a significant role in reducing the total energy cost of phases exhibiting this symmetry (see Sec. III). Moreover, as can be observed in Fig. 1(f), Fig. 3(f), Fig. 5(f), as well as Appendix B, all quasicrystalline phases studied have a sector in which the superfluid fraction is significant, allowing us to conclude that these are actually super quasicrystal phases. These results were confirmed by extensive Gross-Pitaevskii computations, which produce equivalent configuration and superfluid fraction to those obtained with our variational spectral approach.
Although the present work considers a strictly two-dimensional model, the results obtained are also relevant for three-dimensional systems strongly confined along the $z$-direction. In such cases, the confining potential produces a particle distribution $|\psi_{z}(z)|^{2}$, often described by a Gaussian or Thomas–Fermi profile [zhang2019supersolidity, ripley2023two] along the $z$-axis. The total ground-state wave function is then typically written as $\psi_{\mathrm{3D}}(\textbf{x})=\psi_{z}(z)\psi_{\mathrm{2D}}(x,y)$, where $|\psi_{\mathrm{2D}}(x,y)|^{2}$ describes the density pattern of the system in the $xy$-plane. This kind of ansatz enables a dimensional reduction procedure by integrating out the $z$-coordinate in the energy per particle functional, which leads to an effective two-dimensional problem in terms of $\psi_{\mathrm{2D}}(x,y)$. The resulting 22D effective problem has essentially the same structure of the two-dimensional model considered in Eq. (3). In this case, the effective two-dimensional contact interaction corresponds to a “renormalized” 33D contact interaction that depends on the width of the system along the $z$-direction. Furthermore, regarding the laser-induced pair interaction between particles, it can be considered constant along the $z$-direction in the strong confinement regime [bonifacio2024laser]. This implies that the momentum-dependent component of the effective 22D pair interaction will not be affected by the integration over the $z$-coordinate. Therefore, we conclude that all our results concerning the feasibility of producing quasicrystalline phases in two-dimensional bosonic systems can also be extrapolated to the more relevant experimental scenario of three-dimensional systems strongly confined along the $z$-direction.
@@
We believe that the present results are relevant for the investigation of quasicrystal structures produced in cavity mediated interacting Bose-Einstein condensates [mivehvar2019emergent, mivehvar2021cavity, masalaeva2023tuning, orsi2024cavity]. The design on demand of cavity mediated interactions have paved a way to investigate a variety of cutting-edge many-body phases. In this regard, such systems show the interesting opportunity to operate with lattice phononic modes, a flexible manner to test quasicrystals in a quantum framework [mendoza2025excitations]. At the same time, the interplay between long-range magnetic and light-induced interactions in a Bose-Einstein condensate [mishra2023crystalline] is another promising experimental setup that, in principle, should observe 8-QC, 10-QC and 12-QC patterns displaying a finite superfluid signal. Especially relevant is also the usage of dipolar atoms in cavities [bonifacio2024laser], allowing the implementation of many-body systems with highly tunable interactions capable of producing self-assembled quasicrystal clusters. Here, it will be of key importance to inspect the stability of octagonal as well as decagonal phases.
@@ VII Acknowledgments
A.M.C. acknowledge UNIFI for financial support and hospitality. This work was supported by the European Union through the Next Generation EU funds through the Italian MUR National Recovery and Resilience Plan, Mission 4 Component 2–Investment 1.4–National Center for HPC, Big Data and Quantum Computing (CUP Grant No. B83C22002830001). F. C. and V. Z. acknowledge financial support from PNRR Ministero Università e Ricerca Project No. PE0000023-NQSTI funded by European Union-Next-Generation EU. The numerical solution of the GPE was performed using the XMDS2 software [dennis2013xmds2].
Appendix A Decagonal quasicrystal with alternative three-minima pair interaction potential
Figure 7: (a) Pair interaction potential containing a three minima structure in Fourier space. The minima are located at the wave vectors $\tilde{k}_{1}$, $\tilde{k}_{2}$ and $\tilde{k}_{4}$ for the decagonal structure (see Tab. 2). (b) Ground-state phase diagram of the model using $\hat{v}(\bar{k})$, with $\bar{k}=\tilde{k}_{4}$ and $\gamma$ as running parameters. The acronyms HSF, $\mathrm{HS}_{i}$ and 1010-QC stand for homogeneous superfluid, hexagonal solid and decagonal quasicrystal phase, respectively. The sub-index “i” in the HS acronym indicate a phase with characteristic wave vector $\tilde{k}_{i}$ as defined in Tab. 2.
An alternative scenario to the one considered in Fig. 3(e) corresponds to the case in which a third minimum is added at $k=\tilde{k}_{4}$ (see Tab. 2). To make possible the comparison with the phase diagrams in Fig. 3, we consider the situation where $\hat{v}(\tilde{k}_{1})=\hat{v}(\tilde{k}_{2})=-1$. In this way, the resulting pair interaction potential in momentum space, $\hat{v}(k)$, is shown in Fig. 7(a). The corresponding phase diagram, using $\hat{v}(\tilde{k}_{4})$ and $\gamma$ as running parameters, is shown in Fig. 7(b). As can be observed, for sufficiently high values of $\hat{v}(\bar{k})$, the hexagonal phase with characteristic wave vector $\tilde{k}_{1}$ is the most stable modulated phase. As $\hat{v}(\bar{k})$ decreases, the system undergoes a transition to a decagonal quasicrystal ($10$-QC) phase and eventually to an $\mathrm{HS}_{4}$ phase in the low $\hat{v}(\bar{k})$ regime. Additionally, for small values of $\gamma$, the system exhibits a transition to a homogeneous superfluid (HSF) phase, as expected.
By comparing the phase diagram obtained for the present model with that of Fig. 3(e), we observe that the extension of the $10$-QC phase along the $\hat{v}(\bar{k})$ axis is increased in the present case. Moreover, the $10$-QC region in the phase diagram of Fig. 3(e) is shifted toward higher values of $\hat{v}(\bar{k})$. This shift results from the lower momentum of the third minimum excited by the pair potential in that case, which generates density Fourier modes with lower kinetic energy, thus favoring the early stabilization of the $10$-QC phase. Finally, we conclude that, from a general perspective, the study of the model shown in Fig. 7 confirm the necessity of the competition among three length scales to stabilize the $10$-QC phase.
Appendix B Superfluid fraction as a function of $\gamma$
In this appendix, we discuss the behavior of the superfluid fraction as $\gamma$ is varied across various models that display quasicrystalline phases. This study considers three representative examples, one for each type of quasicrystal. Specifically, we compute the superfluid fraction for a fixed pair potential $\hat{v}(k)$, using $\gamma$ as the running parameter. For our analysis, we have selected the models presented in Fig. 1(a), Fig. 3(c), and Fig. 5(c), corresponding to the dodecagonal, decagonal, and octagonal quasicrystals, respectively. As in the main text, two independent methods are employed for this calculation: direct GPE simulations and the spectral variational method.
In Fig. 8, we present the resulting superfluid fractions as a function of $\gamma$, considering $\hat{v}(\tilde{k}_{2})=-2.0$ in panel (a), $\hat{v}(\tilde{k}_{4})=0.0$ in panel (b), and $\hat{v}(\tilde{k}_{4})=-1.0$ in panel (c). As can be observed, within a given modulated phase, the superfluid fraction generally decreases as $\gamma$ increases. This behavior is a direct consequence of the growing modulation amplitude and increased particle localization as the interaction strength, controlled by $\gamma$, becomes stronger. In addition to this monotonic trend, discontinuities in the superfluid fraction are visible at the phase boundaries, as expected in a first-order phase transition scenario.
Figure 8: Superfluid fraction $(f_{s})$ as a function of the pair potential intensity $\gamma$ for the dodecagonal (1212-QC), decagonal (1010-QC) and octagonal (88-QC) quasicrystal structures. (a) Model presented in Fig. 1(a), considering $\hat{v}(\tilde{k}_{2})=-2.0$. (b) Model in Fig. 3(c) for $\hat{v}(\tilde{k}_{4})=0.0$. (c) Model from Fig. 5(c) setting $\hat{v}(\tilde{k}_{4})=-1.0$. Black dashed lines indicate the phase boundaries while the intermediate phase in all diagrams corresponds to the HS1 phase. The blue points corresponds to the results obtained via spectral mean-field method while orange diamonds represent values from GPE simulations.
In the regime of small to moderate values of $\gamma$, the system simultaneously exhibits spatial order and superfluidity in both periodic and aperiodic phases, thereby confirming the existence of supersolid and super quasicrystalline phases. In the large $\gamma$ regime, the superfluid fraction approaches zero; however, it never strictly vanishes. This is a well-known limitation of the mean-field variational description of Bose-Einstein condensates. Finally, we would like to remark the excellent agreement between the results obtained via the two methods employed, confirming the reliability and consistency of the superfluid fraction calculations.
Appendix C Homogeneous–hexagonal phase boundary behavior
The behavior of the homogeneous–hexagonal ($\mathrm{HS}_{1}$) phase boundary can be approximately understood using a Ginzburg–Landau approach in terms of the leading Fourier mode amplitude ($c_{1}$) of the hexagonal pattern. This description is naturally valid in the regime of weak modulations, hopefully occurring in the vicinity of the homogeneous–hexagonal phase boundary.
In this case, the ground state wave function can be written by the following ansatz:
[…]
$+$ $\left.\left.\cos\left(-\frac{kx}{2}-\frac{\sqrt{3}ky}{2}\right)\right)\right],$
where the normalization factor is given as $Z=\sqrt{1+\frac{3c_{1}^{2}}{2}}$. By inserting this ansatz into the energy per particle functional of Eq. (3) and expanding the resulting expression in powers of $c_{1}$, we obtain, up to order $O(c_{1}^{5})$:
$\begin{split}E_{\mathrm{hex}}(c_{1},k)=\,&E_{\mathrm{hom}}+\left(\frac{3\tilde{k}_{1}^{2}}{4}+3\gamma\hat{v}(\tilde{k}_{1})\right)c_{1}^{2}+\alpha_{3}c_{1}^{3}\\ &+\alpha_{4}c_{1}^{4}+\dots,\end{split}$ (10)
where $\alpha_{i}$ are the coefficients of the power-series expansion in $c_{1}$, and $E_{\mathrm{hom}}=\gamma\hat{v}(0)/2$ denotes the energy per particle of the homogeneous phase.
Although the presence of a cubic term indicates the occurrence of a first-order transition, a good estimate of the transition point can be obtained from the instability threshold condition $3\tilde{k}_{1}^{2}/4+3\gamma\hat{v}(\tilde{k}_{1})\leq 0$. This yields an approximate boundary located at
$\gamma=-\frac{\tilde{k}_{1}^{2}}{4\hat{v}(\tilde{k}_{1})}.$ (11)
In all cases considered, the main Fourier component of the $\mathrm{HS}_{1}$ phase corresponds to $\tilde{k}_{1}=1$, while the value of $\hat{v}(\tilde{k}_{1})$ is fixed to 1-1. Thus, the position of the phase boundary is estimated as $\gamma=0.25$ for all cases. This observation explains why the position of the homogeneous to $\mathrm{HS}_{1}$ transition is essentially constant for all phase diagrams. Moreover, it is important to mention that the position for the phase boundary evaluated here is an estimate based on the leading Fourier harmonic expansion for the hexagonal pattern. The actual value observed in the phase diagrams is slightly lower due to the relevant role of higher order Fourier harmonics in the wave function expansion.